# Relational Quantum Mechanics with Cross-Perspective Links Postulate: an Internally Inconsistent Scheme

### 12 Dec 2023 v1

@@ Abstract
We discuss the status of relative facts – the central concept of Relational Quantum Mechanics (RQM) – in the context of the new amendment to RQM called cross-perspective links postulate. The new axiom states that by a proper measurement one learns the value of the relative outcome/fact earlier obtained by another observer-system. We discuss a Wigner-Friend-type scenario in which, without cross-perspective links postulate, relative facts have no any predictive or causal power, whereas including cross-perspective links makes them effectively hidden variables, which causally determine outcomes of specific measurements. However, cross-perspective links axiom invalidates the other axiom of RQM, the one which states that in a Wigner-Friend scenario, RQM assigns an entangled state to the Friend and System after the unitary transformation of their interaction, despite the appearance of the relative fact for the Friend. This quantum mechanical state according to RQM properly describes the situation for Wigner. This shows that RQM with cross-perspective links axiom is an internally inconsistent hidden variable theory and therefore cannot be treated as an interpretation of quantum mechanics in any sense.
@@ I Introduction
  • “cross-perspective links:” In a scenario where some observer Alice measures a variable $V$ of a system $S$, then provided that Alice does not undergo any interactions which destroy the information about $V$ stored in Alice’s physical variables, if Bob subsequently measures the physical variable representing Alice’s information about the variable $V$, then Bob’s measurement result will match Alice’s measurement result.  [^1] The cross-perspective-links axiom is in [1] numbered 4.1, as it is to replace axiom RQM-4. We will be referring to this axiom throughout the paper either as cross-perspective links or RQM-4.1
@@
One should comment here that the word "realisation" should, we think, be replaced here by "RQM main hypothesis". We also cannot agree, that the extension is "minimal". Still, this quotation defines the RQM standpoint.
@@ I.2 The codification of RQM, and its change
  • RQM-5. Measurement: an interaction between two systems results in a correlation within the interactions between these two systems and a third one; i.e. with respect to a third system $W$, the interaction between the two systems $S$ and $F$ is described by a unitary evolution that potentially entangles the quantum states of $S$ and $F$.
  • RQM-6. Internally consistent descriptions: In a scenario where $F$ measures $S$, and $W$ also measures $S$ in the same basis, and $W$ then interacts with $F$ to ‘check the reading’ of a pointer variable (i.e. by measuring $F$ in the appropriate ‘pointer basis’), the two values found are in agreement.
@@ Appendix A Observables of compound quantum systems
does not necessarily hold in RQM, since [[3], p. 4]:

### 03 Jan 2024 v2

@@ Abstract
We discuss the status of relative facts – the central concept of Relational Quantum Mechanics (RQM) – in the context of the new amendment to RQM called cross-perspective links postulate. The new axiom states that by a proper measurement one learns the value of the relative outcome/fact earlier obtained by another observer-system. We discuss a Wigner-Friend-type scenario in which, without cross-perspective links postulate, relative facts have no predictive or causal power, whereas including cross-perspective links makes them effectively hidden variables, which causally determine outcomes of specific measurements. However, cross-perspective links axiom invalidates the other axiom of RQM, the one which states that in a Wigner-Friend scenario, RQM assigns an entangled state to the Friend and System after the unitary transformation of their interaction, despite the appearance of the relative fact for the Friend. This quantum mechanical state according to RQM properly describes the situation for Wigner. From this we show that RQM with cross-perspective links axiom is an internally inconsistent hidden variable theory and therefore cannot be treated as an interpretation of quantum mechanics in any sense.
@@ I Introduction
  • “cross-perspective links:” In a scenario where some observer Alice measures a variable $V$ of a system $S$, then provided that Alice does not undergo any interactions which destroy the information about $V$ stored in Alice’s physical variables, if Bob subsequently measures the physical variable representing Alice’s information about the variable $V$, then Bob’s measurement result will match Alice’s measurement result.  [^1] The cross-perspective-links axiom is in the preprint version of [1] numbered 4.1, as it is to replace axiom RQM-4. We will be referring to this axiom throughout the paper either as cross-perspective links or RQM-4.1
@@
One should comment here that the word "realisation" should, we think, be replaced here by "RQM assertion". We also cannot agree, that the extension is "minimal". Still, this quotation defines the RQM standpoint.
@@ I.2 The codification of RQM, and its change
  • RQM-5. Measurement: an interaction between two systems results in a correlation within the interactions between these two systems and a third one; that is, with respect to a third system $W$, the interaction between the two systems $S$ and $F$ is described by a unitary evolution that potentially entangles the quantum states of $S$ and $F$.
  • RQM-6. Internally consistent descriptions: In a scenario where $F$ measures $S$, and $W$ also measures $S$ in the same basis, and $W$ then interacts with $F$ to ‘check the reading’ of a pointer variable (i.e., by measuring $F$ in the appropriate ‘pointer basis’), the two values found are in agreement.
@@ Appendix A Observables of compound quantum systems
does not necessarily hold in RQM, since [[3], p. 5]: